Show your complete solution.


1. Use Lagrange Multiplier to determine the dimensions of a rectangular box, open at the top, having a volume of 32 cubic feet and requiring the least amount of material for its construction.


1
Expert's answer
2020-04-20T15:20:49-0400

Let dimensions of the box be a, b and c feet respectively.

Rectangular box, open at the op

Given : "Volume = abc=32ft^3"

To find :  least amount of material for its construction, which means minimum surface area.

Solution : "S=ab+2bc+2ca" is the surface area of the required box.


Thus, "V(a,b,c)=abc-32=0"

And "S(a,b,c)=ab+2bc+2ca" = minimum.

Using the Method of Lagrange's Multiplier, we get;

"\triangledown S=\lambda \triangledown V"

"\triangledown S=\begin{bmatrix}
 S_a\\ S_b \\S_c
\end{bmatrix}=\begin{bmatrix}
 b+2c\\ a+2c \\2a+2b
\end{bmatrix};" "\triangledown V=\begin{bmatrix}
 V_a\\ V_b \\V_c
\end{bmatrix}=\begin{bmatrix}
 bc\\ ac \\ab
\end{bmatrix}"

"\implies \begin{bmatrix}
 b+2c\\ a+2c \\2a+2b
\end{bmatrix}=\lambda \begin{bmatrix}
 bc\\ ac \\ab
\end{bmatrix}"

"\implies \begin{bmatrix}
 b+2c\\ a+2c \\2a+2b
\end{bmatrix}= \begin{bmatrix}
 \lambda bc\\ \lambda ac \\\lambda ab
\end{bmatrix}"

"\implies \lambda = (b+2c)/bc=(a+2c)/ac=(2a+2b)/ab"

"(b+2c)/bc=(a+2c)/ac \implies 1+2c/b=1+2c/a"

"\implies b=a"

"(a+2c)/ac=(2a+2b)/ab \implies (a+2c)/ac=""4a/a^2=4/a \implies a/c+2=4"

"\implies c=a/2"


"V=abc-32=a(a)(a/2)-32=0 \implies a^3-64 =0"

"\implies a=4; b=4; c=2"

"\implies S= ab+2bc+2ca=16+16+16=48 ft^2"


Thus,48 square feet is the least amount of material for its construction.